The finished inside diameter of a piston ring is normally distributed with a mean of 10 cm and a standard deviation of 0.03 cm. Below what value of inside diameter will 15% of the piston rings fall?


1
Expert's answer
2021-05-17T17:30:01-0400

Z multiplier corresponding to probability of 0.15 is -1.036

Z=(X-mu)/sigma=-1.036

(X-10)/0.03=-1.036

X=-1.036*0.03+10

X=9.9689